# Watch and Listen: Understanding Audio-Visual-Speech Moments with Multimodal LLM

### v2

@@ Abstract
Humans naturally understand moments in a video by integrating visual and auditory cues. For example, localizing a scene in the video like “A scientist passionately speaks on wildlife conservation as dramatic orchestral music plays, with the audience nodding and applauding” requires simultaneous processing of visual, audio, and speech signals. However, existing models often struggle to effectively fuse and interpret audio information, limiting their capacity for comprehensive video temporal understanding. To address this, we present TriSense, a triple-modality large language model designed for holistic video temporal understanding through the integration of visual, audio, and speech modalities. Central to TriSense is a Query-Based Connector that adaptively reweights modality contributions based on the input query, enabling robust performance under modality dropout and allowing flexible combinations of available inputs. To support TriSense’s multimodal capabilities, we introduce TriSense-2M, a high-quality dataset of over 2 million curated samples generated via an automated pipeline powered by fine-tuned LLMs. TriSense-2M includes long-form videos and diverse modality combinations, facilitating broad generalization. Extensive experiments across multiple benchmarks demonstrate the effectiveness of TriSense and its potential to advance multimodal video analysis. Code and dataset will be publicly released.
@@ 4.1 Multimodal Information Extraction
During training, given a video $V=\left\{F_{1},F_{2},....F_{T}\right\}$, where $T$ denotes the total number of frames, we first uniformly select $n=64$ frames for further processing and controlling memory usage. For each selected frame, we record its timestamp and extract an audio segment of ±$\pm 1$ second around the frame, resulting in a sequence of audio segments $A=\left\{A_{1},A_{2},....A_{n}\right\}$, each lasting 2 seconds. For instance, if a frame is sampled at 123.4 seconds, the corresponding audio segment spans from 122.4 to 124.4 seconds. We then use pretrained expert encoders to extract modality-specific tokens: $f^{v}_{i}$, $f^{a}_{i}$, $f^{s}_{i}$ for vision, audio, and speech, repectively.
@@ 4.2 Query-Based Connector
To more effectively integrate multimodal inputs with the query and enhance sensitivity to salient features, we introduce a Query-Based Connector that adaptively balances the contributions of each modality based on the query’s content, as illustrated in Figure 4. Detailed pseudocode is provided in the Appendix. The compressed modality features $f^{{}^{\prime}v}_{i}$, $f^{{}^{\prime}a}_{i}$ and $f^{{}^{\prime}s}_{i}$ obtained from Section 4.1, are passed through Cross-Attention layers, where they interact with the encoded query representation $f^{(q)}$. The objective is for each attention layer to emphasize features that are most relevant to the query. The outputs of these layers are denoted as query-relevant features $f^{v,q}_{i}$, $f^{a,q}_{i}$ and $f^{s,q}_{i}$, which reflect the alignment between each modality and the query.
@@
[…]
@@ 4.3 Causal Event Prediction
To support both temporal and textual outputs, we introduce adaptive head switching via a special 〈sync〉 token during LLM generation. This token is appended to the vocabulary and serves as a control signal that guides the model to switch between time head and language model (LM) head, as illustrated in Figure  4. When the 〈sync〉 token is encountered, the LLM transitions between decoding modalities to generate either timestamp-aligned predictions or free-form textual outputs, depending on the task.
@@ B M R C B M R C B M R C B M R C
5.1 Evaluation Datasets, Metrics and Baseline Models
  • [label=•,leftmargin=12pt,nosep]
  • Moment Retrieval (MR). In this task, the model is required to retrieve specific segments within a video that correspond to a given textual query. We evaluate performance on TriSense-2M, as well as two widely used public benchmarks: Charades-ST [8] and ActivityNet-Captions [13]. Retrieval effectiveness is reported using Recall@IoU=0.5, Recall@IoU=0.7, and mean IoU (mIoU), providing a comprehensive view of the model’s localization accuracy at varying overlap thresholds.
@@ 5.1 Evaluation Datasets, Metrics and Baseline Models
The modality combinations are set to Audio-Visual-Speech (AVS), Visual-Speech (VS), Audio-Visual (AV) and Visual-Only (V). To establish a solid comparative baseline, we select representative models specifically designed for Video Temporal Grounding (VTG) tasks. These include VTimeLLM [15], TimeChat [27], VTG-LLM [12], TRACE [11], along with two recent omni-modal models LongVALE [10] and Qwen2.5-Omni [34]. While Momentor [23], Hawkeye [31] and NumPro-FT [33] are not included in the TriSense-2M benchmark due to the unavailability of model checkpoints and their lack of support for captioning, they are considered in public benchmark evaluations as reported in their official papers. By comparing our model against these baselines across diverse tasks and evaluation metrics, we aim to provide a comprehensive assessment of TriSense’s capabilities and its advancements in video temporal understanding.
5.2 Results and Analysis
Superior Performance on Omni-Modal datasets. We evaluate our performance on the proposed TriSense-2M dataset. As illustrated in Table 2 and Table 3, TriSense consistently outperforms existing video LLMs across nearly all evaluated tasks. It also significantly surpasses latest omni-modal models like LongVALE [10] and Qwen2.5-Omni [34], particularly in the audio-visual-speech (AVS) setting where all three modalities are leveraged. Due to page limitations, additional zero-shot performance results on the LongVALE benchmark are provided in the Appendix.
@@ 5.2 Results and Analysis
Zero-shot performance on public Moment Retrieval benchmarks. We evaluate TriSense in a zero-shot setting on established visual-only datasets, including Charades-STA [8] and ActivityNet-Captions [13], as shown in Table 4. The results show that although TriSense shows slight inferior performance in Table 3, it still achieves competitive performance in visual-only settings, showing especially higher accuracy (IoU=0.7) than others, even with less frames used.
@@
Ablation Studies. We conduct ablation experiments to assess the contribution of different components, including the training strategy, the Query-Based Connector, and the number of frames processed by the model. Due to page limit, we include only the moment retrieval ablation results in the main text, while segment captioning ablations are presented in the Appendix. As shown in Table 5, we compare our adaptive weighting strategy against simpler alternatives. The Addition baseline directly sums the modality features without weighting. The Fixed Weights baseline assigns equal weights (e.g., 0.33 each in AVS), or fixed values depending on the modality pair (e.g., 0.5 for each active modality and 0 for inactive ones in VS/AV), and uses only the visual stream in visual-only tasks (weight of 1, others set to 0). These comparisons confirm the effectiveness of our query-adaptive weighting mechanism.
We can observe that in Training Stages section, Stage 1 focuses solely on modality alignment and no temporal information is included, therefore does not performing well on moment retrieval. However, after training in Stage 2, the model acquires around 50% of its moment retrieval capability. For the Connector ablation, we find that simply adding all modalities together does not allow the model to emphasize the more important modalities, resulting in a drop in performance. Similarly, in the Fixed Weights ablation for AVS/VS/AV tasks, assigning equal weights to modalities fails to effectively capture their varying importance, leading to inferior performance compared to using dynamic modality weights. However, fixing visual weight to 1 indeed leads to slightly better performance in visual-only (V) setting. We also observe that increasing the number of frames used by the model (from 64 to 128) leads to performance improvement in every scenario. This trend is also consistent with the findings in Segment Captioning presented in the Appendix.
@@ A.1 Training Recipe
Our training process follows a structured three-stage approach: Feature Alignment, Connector Generalization, and Instruction Tuningg—to progressively equip the model with strong multimodal and temporal reasoning capabilities. The specific components trained at each stage are illustrated in Figure 4.
@@ A.2 Datasets
To avoid massive evaluation time, we extract 11,415 challenging samples from the 0.1M test set using two filtering criteria: 1) The majority of events should occur in the middle portion of the video rather than the beginning. 2) Captions must contain at least 20 words. Evaluation is conducted using a single A100 SXM4 80GB GPU with a batch size of 1, requiring approximately 8–10 hours to complete. All models in comparison are evaluated on this same test subset, using their officially recommended hyperparameters (e.g., number of frames, temperature, top-p, etc.)
@@ A.3 Detailed training settings
Our multimodal framework incorporates dedicated encoders for each modality. For the visual modality, we adopt openai/clip-vit-large-patch14-336 [25]; for audio and speech modalities, we employ BEATs_iter3+ (AS2M) (cpt2) [4] and Whisper-large-V3 [26], respectively. As for the large language model (LLM) backbone, we select Mistral-7B [17], initialized from TRACE [11], instead of using other LLM backbones. , rather than using other LLM backbones. This choice is motivated by TRACE’s prior training on large-scale temporal understanding data, which equips it with stronger temporal reasoning capabilities. The maximum context length is configured to 4096 tokens.
@@
Training Duration 10 Hours 3.5 days & 7 days
@@
In Stage 1, we train the model with a batch size of 512 and single-frame input, completing within 10 hours using 4$\times$×A100SXM4-80GB GPUs. Stage 2 and 3 are conducted on 16$\times$×A100SXM4-80GB GPUs with batch sizes of 128 and 256, respecitvely. Stage 2 requires approximately 3.5 days to complete, and stage 3 requires 7 days to finish. We use DeepSpeed Zero2 because the BEATs model does not function properly under Zero3 settings. Further details on datasets and hyperparameters are provided in 7.
Appendix B Data Format
@@ B.2 Training data format
Following the approach in TRACE, we use special tokens such as 〈sync〉 and 〈time〉 to signal the model to switch between different prediction heads. An example of the data structure is illustrated in 6.
@@
C.1 Ablation study on Segment Captioning
We present ablation results for the Segment Captioning task in Table 8. In the Training Stages section, we observe that even though Stage 1 does not involve temporal modeling, the model still demonstrates some effectiveness on segment captioning. This is attributed to the modality-text alignment enabled by paired training, which facilitates basic understanding across modalities. After Stage 2, where the training incorporates the Query-Based Connector, the temporal module, and the LM head (with the LLM backbone still frozen), the model reaches about 70–80% of the full model’s performance, indicating the benefit of temporal component training.
In the Connector ablation, removing adaptive weighting leads to degraded performance. The Addition variant and the first three configurations of the Fixed Weights setup show weak results, with Addition performing the worst. However, similar to findings in the Moment Retrieval task, assigning a fixed weight of 1 to visual tokens in visual-only settings produces a minor performance boost.
In the Frame Number analysis, increasing the number of input frames yields moderate gains in most cases, echoing the trend seen in Section 5.2. However, this improvement is relatively limited. A potential explanation lies in the inherent differences between Segment Captioning and Moment Retrieval. While both require temporal reasoning, Segment Captioning relies more on spatial information, and increasing frames may introduce spatial redundancy that clashes with the model’s 64-frame training setup. Conversely, Moment Retrieval benefits more from longer temporal context, where additional frames enrich sequential understanding, resulting in more substantial performance improvements.
Model Frame Number AVS-SC VS-SC AV-SC V-SC
B M R C B M R C B M R C B M R C
Training Stages
Connector
Frame Number
C.2 Omni-modal dataset & General understanding dataset
We conduct additional experiments on the public Omni-Modal benchmark LongVALE [10]. LongVALE is designed for event understanding across vision, audio, and language modalities, comprising 105,000 omni-modal events with precise temporal annotations and relation-aware captions, collected from 8,400 high-quality long-form videos. The Omni-VTG and Omni-SC tasks named from the official report of LongVALE, which are the same as AVS-MR and AVS-SC in our paper. As summarized in Table 9, our zero-shot performance on the Moment Retrieval task is comparable to LongVALE’s performance, even though their model is trained on the same dataset. Although there is a larger gap in the Segment Captioning task, we believe this is due to pattern differences in captioning styles between our SC training data and LongVALE’s SC data. Such differences in caption patterns can lead to noticeable drops across all four evaluation metrics.
Table 9: Performance on public Omni-Modal benchmark LongVALE [10]. "*" indicates this model is trained on the LongVALE dataset. ZS and FT represent zero-shot and fine-tuned, respectively. The best and second results are highlighted in bold and underlined, respectively.
Model Omni-VTG (AVS-MR) Omni-SC (AVS-SC)
R@0.3 R@0.5 R@0.7 mIoU B M R C
VideoChat (ZS) 2.2 0.9 0.4 3.0 0.5 9.6 0.0 8.2
VideoChatGPT (ZS) 4.9 2.9 0.9 5.0 0.4 14.0 0.9 5.9
VideoLLaMA (ZS) 2.5 1.1 0.3 1.9 0.9 11.5 0.1 8.9
PandaGPT (ZS) 2.5 1.0 0.3 2.2 0.6 14.9 0.3 8.9
NExT-GPT (ZS) 4.3 1.9 0.7 4.0 0.4 10.2 0.0 8.1
TimeChat (ZS) 5.8 2.6 1.1 5.2 1.2 16.1 1.6 10.0
VTimeLLM (ZS) 7.5 3.4 1.3 6.4 1.0 14.5 1.6 5.5
LongVALE (FT) * 15.7 8.6 3.9 11.0 5.6 22.4 20.3 10.9
TriSense (ZS) 14.8 9.3 4.7 11.2 4.8 21.9 18.8 10.4

### v3

@@ Abstract
Humans naturally understand moments in a video by integrating visual and auditory cues. For example, localizing a scene in the video like “A scientist passionately speaks on wildlife conservation as dramatic orchestral music plays, with the audience nodding and applauding” requires simultaneous processing of visual, audio, and speech signals. However, existing models often struggle to effectively fuse and interpret audio information, limiting their capacity for comprehensive video temporal understanding. To address this, we present TriSense, a triple-modality large language model designed for holistic video temporal understanding through the integration of visual, audio, and speech modalities. Central to TriSense is a Query-Based Connector that adaptively reweights modality contributions based on the input query, enabling robust performance under modality dropout and allowing flexible combinations of available inputs. To support TriSense’s multimodal capabilities, we introduce TriSense-2M, a high-quality dataset of over 2 million curated samples generated via an automated pipeline powered by fine-tuned LLMs. TriSense-2M includes long-form videos and diverse modality combinations, facilitating broad generalization. Extensive experiments across multiple benchmarks demonstrate the effectiveness of TriSense and its potential to advance multimodal video analysis.
@@ 3 Data Construction
Table 2: Detailed Information of TriSense-2M, where Avg/Min/Max Duration represent the average, minimum, and maximum duration, respectively. 0–5s / 5–10s / 10–15s, etc., represent the proportions of different duration intervals.
Total Events Avg Duration Min Duration Max Duration 0-5s 5-10s 10-15s 15-20s 20-30s
1940522 6.87s 2.00s 30.00s 35.5% 46.2% 12.6% 4.2% 1.5%
@@ 4.1 Multimodal Information Extraction
During training, given a video $V=\left\{F_{1},F_{2},....F_{T}\right\}$, where $T$ denotes the total number of frames, we first uniformly select $n=64$ frames for further processing and controlling memory usage. For each selected frame, we record its timestamp and extract an audio segment of $\pm 1$ second around the frame, resulting in a sequence of audio segments $A=\left\{A_{1},A_{2},....A_{n}\right\}$, each lasting 2 seconds. For instance, if a frame is sampled at 123.4 seconds, the corresponding audio segment spans from 122.4 to 124.4 seconds. We then use pretrained expert encoders to extract modality-specific tokens: $f^{v}_{i}$, $f^{a}_{i}$, $f^{s}_{i}$ for vision, audio, and speech, repectively.
@@ 4.2 Query-Based Connector
To more effectively integrate multimodal inputs with the query and enhance sensitivity to salient features, we introduce a Query-Based Connector that adaptively balances the contributions of each modality based on the query’s content, as illustrated in Figure 4. The compressed modality features $f^{{}^{\prime}v}_{i}$, $f^{{}^{\prime}a}_{i}$ and $f^{{}^{\prime}s}_{i}$ obtained from Section 4.1, are passed through Cross-Attention layers, where they interact with the encoded query representation $f^{(q)}$. The objective is for each attention layer to emphasize features that are most relevant to the query. The outputs of these layers are denoted as query-relevant features $f^{v,q}_{i}$, $f^{a,q}_{i}$ and $f^{s,q}_{i}$, which reflect the alignment between each modality and the query.
@@
Here, $\mathbf{c}_{m}$ is the compressed modality vector after global average pooling, $T_{m}$} indicates the sequence length of distinct modality $m$, $\|$ denotes vector concatenation, and $\tilde{w}_{m}$ and $w_{m}$ represent unnormalized and normalized weights, respectively.
@@ 4.3 Causal Event Prediction
To support both temporal and textual outputs, we introduce adaptive head switching via a special ⟨sync⟩ token during LLM generation. This token is appended to the vocabulary and serves as a control signal that guides the model to switch between time head and language model (LM) head, as illustrated in Figure  4. When the ⟨sync⟩ token is encountered, the LLM transitions between decoding modalities to generate either timestamp-aligned predictions or free-form textual outputs, depending on the task.
@@ 5 Experiments
5.1 Evaluation Datasets, Metrics and Baseline Models
  • [label=•,leftmargin=12pt,nosep]
  • Moment Retrieval (MR). In this task, the model is required to retrieve specific segments within a video that correspond to a given textual query. We evaluate performance on TriSense-2M, as well as two widely used public benchmarks: Charades-STA [8] and ActivityNet-Captions [13]. Retrieval effectiveness is reported using Recall@IoU=0.5, Recall@IoU=0.7, and mean IoU (mIoU), providing a comprehensive view of the model’s localization accuracy at varying overlap thresholds.
The modality combinations are set to Audio-Visual-Speech (AVS), Visual-Speech (VS), Audio-Visual (AV) and Visual-Only (V). To establish a solid comparative baseline, we select representative models specifically designed for Video Temporal Grounding (VTG) tasks. These include VTimeLLM [15], TimeChat [27], VTG-LLM [12], TRACE [11], along with two recent omni-modal models LongVALE [10] and Qwen2.5-Omni [34]. While Momentor [23], Hawkeye [31] and NumPro-FT [33] are not included in the TriSense-2M benchmark due to the unavailability of model checkpoints and their lack of support for captioning, they are included in public benchmark evaluations according to the reports in their official papers. By comparing our model against these baselines across diverse tasks and evaluation metrics, we aim to provide a comprehensive assessment of TriSense’s capabilities and its advancements in video temporal understanding.
5.2 Results and Analysis
Superior Performance on Omni-Modal datasets. We evaluate our performance on the proposed TriSense-2M dataset. As illustrated in Table 3 and Table 4, TriSense consistently outperforms existing video LLMs across nearly all evaluated tasks. It also significantly surpasses latest omni-modal models like LongVALE [10] and Qwen2.5-Omni [34], particularly in the audio-visual-speech (AVS) setting where all three modalities are leveraged.
@@ B M R C B M R C B M R C B M R C
We also conduct experiments on the public Omni-Modal benchmark LongVALE [10]. LongVALE is designed for event understanding across vision, audio, and language modalities, comprising 105,000 omni-modal events with precise temporal annotations and relation-aware captions, collected from 8,400 high-quality long-form videos. The Omni-VTG and Omni-SC tasks named from the official report of LongVALE, which are the same as AVS-MR and AVS-SC in our paper. As summarized in Table 5, our zero-shot performance on the Moment Retrieval task is comparable to LongVALE’s performance, even though their model is trained on the same dataset. Although there is a larger gap in the Segment Captioning task, we believe this is due to pattern differences in captioning styles between our Segment Captioning training data and LongVALE’s Segment Captioning data. Such differences in caption patterns can lead to noticeable drops across all four evaluation metrics.
Table 5: Performance on public Omni-Modal benchmark LongVALE [10]. "*" indicates this model is trained on the LongVALE dataset. The best and second results are highlighted in bold and underlined, respectively.
Model Omni-VTG (AVS-MR) Omni-SC (AVS-SC)
R@0.3 R@0.5 R@0.7 mIoU B M R C
VideoChat (7B) 2.2 0.9 0.4 3.0 0.5 9.6 0.0 8.2
VideoChatGPT (7B) 4.9 2.9 0.9 5.0 0.4 14.0 0.9 5.9
VideoLLaMA (7B) 2.5 1.1 0.3 1.9 0.9 11.5 0.1 8.9
PandaGPT (7B) 2.5 1.0 0.3 2.2 0.6 14.9 0.3 8.9
NExT-GPT (7B) 4.3 1.9 0.7 4.0 0.4 10.2 0.0 8.1
TimeChat (7B) 5.8 2.6 1.1 5.2 1.2 16.1 1.6 10.0
VTimeLLM (7B) 7.5 3.4 1.3 6.4 1.0 14.5 1.6 5.5
LongVALE* (7B) 15.7 8.6 3.9 11.0 5.6 22.4 20.3 10.9
TriSense (7B) 14.8 9.3 4.7 11.2 4.8 21.9 18.8 10.4
@@
Zero-shot performance on public Moment Retrieval benchmarks. We also evaluate TriSense in a zero-shot setting on established two classical visual-only datasets, including Charades-STA [8] and ActivityNet-Captions [13], as shown in Table 6. The results show that although TriSense shows slight inferior performance in Table 4, it still achieves competitive performance in visual-only settings, showing especially higher accuracy (IoU=0.7) than others, even with less frames used.
@@
Ablation Studies. We conduct ablation experiments to assess the contribution of different components, including the training strategy, the Query-Based Connector, and the number of frames processed by the model. As shown in Table 7 and Table 8, we compare our adaptive weighting strategy against simpler alternatives. The Addition baseline directly sums the modality features without weighting. The Fixed Weights baseline assigns equal weights (e.g., 0.33 each in AVS), or fixed values depending on the modality pair (e.g., 0.5 for each active modality and 0 for inactive ones in VS/AV), and uses only the visual stream in visual-only tasks (weight of 1, others set to 0). These comparisons confirm the effectiveness of our query-adaptive weighting mechanism.
Model Frame Number AVS-SC VS-SC AV-SC V-SC
B M R C B M R C B M R C B M R C
Training Stages
Connector
Frame Number
We can observe that in Training Stages section, Stage 1 focuses solely on modality alignment and no temporal information is included, therefore does not performing well across two tasks. However, after training in Stage 2, the model acquires around 50% of its capability. For the Connector ablation, we find that simply adding all modalities together does not allow the model to emphasize the more important modalities, resulting in a drop in performance. Similarly, in the Fixed Weights ablation for AVS/VS/AV tasks, assigning equal weights to modalities fails to effectively capture their varying importance, leading to inferior performance compared to using dynamic modality weights. However, fixing visual weight to 1 indeed leads to slightly better performance in visual-only (V) setting. We also observe that increasing the number of frames used by the model (from 64 to 128) leads to performance improvement in every scenario. This trend is consistent with the findings in both of the Moment Retrieval and Segment Captioning.
@@ 6 Conclusion
7 Acknowledgements
This project is jointly supported by the University of Western Australia (UWA) HDR Scholarship, Australian Research Council ARC DP210101682, DP210102674, and the Australian Government through the Australian Research Council’s DECRA funding scheme DE250100030. We thank Zhejiang Lab for providing the computational resources.
@@ A.1 Training Recipe
Our training process follows a structured three-stage approach: Feature Alignment, Connector Generalization, and Instruction Tuning—to progressively equip the model with strong multimodal and temporal reasoning capabilities. The specific components trained at each stage are illustrated in Figure 4.
@@ A.2 Datasets
Table 10: Details of test set.
Num. Videos Num. Events Avg. Event Duration Total Tasks
3805 11415 7.23s 91320
To avoid massive evaluation time, we extract 11,415 challenging samples from the 0.1M test set, as shown in Table˜10 using two filtering criteria: 1) The majority of events should occur in the middle portion of the video rather than the beginning. 2) Captions must contain at least 20 words. Evaluation is conducted using a single A100 SXM4 80GB GPU with a batch size of 1, requiring approximately 8–10 hours to complete. All models in comparison are evaluated on this same test subset, using their officially recommended hyperparameters (e.g., number of frames, temperature, top-p, etc.)
@@ A.3 Detailed training settings
Our multimodal framework incorporates dedicated encoders for each modality. For the visual modality, we adopt openai/clip-vit-large-patch14-336 [25]; for audio and speech modalities, we employ BEATs_iter3+ (AS2M) (cpt2) [2] and Whisper-large-V3 [26], respectively. As for the large language model (LLM) backbone, we select Mistral-7B [17], initialized from TRACE [11], instead of using other LLM backbones, rather than using other LLM backbones. This choice is motivated by TRACE’s prior training on large-scale temporal understanding data, which equips it with stronger temporal reasoning capabilities. The maximum context length is configured to 4096 tokens.
@@
Training Duration 10 Hours 3.5 days & 5.5 days
@@
[…]
Appendix B More details of TriSense-2M
@@ B.2 Training data format
Following the approach in TRACE, we use special tokens such as ⟨sync⟩ and ⟨time⟩ to signal the model to switch between different prediction heads. An example of the data structure is illustrated in 6.
@@
C.1 More ablation studies
We conduct small-scale ablation studies on slot compression and task-specific heads. Specifically, we randomly selected 20K samples and trained for 3 epochs, initializing all weights from [11]. All experiments were conducted with a sampled frame rate of 64. In terms of token compression, we explored different slot configurations, including 8, 32, and 64. For the prediction head, we removed the Time Head and instead encoded temporal information into special tokens, which were injected into the LLM. Prediction was then performed using only the LM Head.
Table 12: Ablation study on the slot compression [12].
AVS-SC AVS-MR Avg. Event Duration
Slot=8 B: 1.6 M: 3.1 R: 13.1 C: 1.9 IoU=0.5: 0.3 IoU=0.7: 0.1
Slot=16 B: 1.8 M: 3.4 R: 14.6 C: 2.4 IoU=0.5: 0.4 IoU=0.7: 0.3
Slot=32 B: 1.8 M: 3.4 R: 15.6 C: 2.4 IoU=0.5: 0.4 IoU=0.7: 0.3
Slot=64 B: 1.9 M: 3.5 R: 16.1 C: 2.4 IoU=0.5: 0.5 IoU=0.7: 0.2
Table 13: Ablation study on the task-specific heads.
AVS-SC AVS-MR
Unified Head B: 1.5 M: 3.1 R: 13.2 C: 1.1 IoU=0.5: 0.2 IoU=0.7: 0.1
Time Head + LM Head B: 1.8 M: 3.4 R: 14.6 C: 2.4 IoU=0.5: 0.4 IoU=0.7: 0.3
As shown in Table˜12, for slot compression, although reducing the compression rate (by increasing the number of slots) can bring slight performance improvements, it also results in a significant increase in computational overhead. Therefore, the benefits are subject to diminishing returns. As shown in the table, setting the slot to 16 currently offers a good balance between performance and computational cost. For task-specific head in Table˜13, using a single unified head alone leads to a significant drop in performance, indicating that the time head provides additional temporal information which is beneficial to the model. There may be better unified solutions in the future.
Table 14: Ablation studies on the branches of Query-based Connector on Segment Captioning tasks.
Model AVS-SC VS-SC AV-SC V-SC
V-Only B: 1.0 M: 1.7 R: 10.1 C: 0.9 B: 1.1 M: 1.6 R: 10.3 C: 1.1 B: 1.0 M: 1.7 R: 10.2 C: 1.1 B: 1.0 M: 1.8 R: 13.3 C: 1.4
VS-Only B: 1.7 M: 2.9 R: 13.3 C: 2.4 B: 1.6 M: 2.9 R: 14.7 C: 2.5 B: 0.7 M: 2.3 R: 13.2 C: 1.5 B: 0.7 M: 1.6 R: 13.2 C: 1.2
AV-Only B: 1.7 M: 3.1 R: 13.7 C: 1.9 B: 1.4 M: 2.2 R: 12.0 C: 2.1 B: 0.9 M: 3.1 R: 15.0 C: 1.7 B: 0.7 M: 1.6 R: 13.0 C: 1.1
AVS-Full B: 1.8 M: 3.4 R: 14.6 C: 2.4 B: 1.7 M: 3.1 R: 15.1 C: 2.5 B: 1.0 M: 3.2 R: 15.3 C: 1.7 B: 0.8 M: 1.8 R: 13.4 C: 1.5
Table 15: Ablation studies on the branches of Query-based Connector on Moment Retrieval tasks.
Model AVS-MR VS-MR AV-MR V-MR
V-Only IoU=0.5: 0.1 IoU=0.7: 0.0 IoU=0.5: 0.1 IoU=0.7: 0.0 IoU=0.5: 0.1 IoU=0.7: 0.1 IoU=0.5: 0.4 IoU=0.7: 0.2
VS-Only IoU=0.5: 0.4 IoU=0.7: 0.1 IoU=0.5: 0.3 IoU=0.7: 0.2 IoU=0.5: 0.1 IoU=0.7: 0.0 IoU=0.5: 0.3 IoU=0.7: 0.1
AV-Only IoU=0.5: 0.4 IoU=0.7: 0.2 IoU=0.5: 0.1 IoU=0.7: 0.1 IoU=0.5: 0.3 IoU=0.7: 0.1 IoU=0.5: 0.3 IoU=0.7: 0.1
AVS-Full IoU=0.5: 0.4 IoU=0.7: 0.3 IoU=0.5: 0.3 IoU=0.7: 0.2 IoU=0.5: 0.3 IoU=0.7: 0.2 IoU=0.5: 0.4 IoU=0.7: 0.2
We also provide more ablation studies in Table˜14 and Table˜15 for different branches of the Query-Based Connector based on the same 20K randomly selected samples as above. For example, AV-Only in the table means we discarded the Speech branch. During training, all the weigths are initialized from [11], and LLM backbone is set to frozen. Results show that all three modalities contribute positively to all tasks, with the full AVS configuration performing best.
C.2 General understanding dataset